# Efficient simulation of the SABR model

### v1

@@ Abstract
We propose an efficient and reliable simulation scheme for the stochastic-alpha-beta-rho (SABR) model. The two challenges of the SABR simulation lie in sampling (i) the integrated variance conditional on terminal volatility and (ii) the terminal price conditional on terminal volatility and integrated variance. For the first sampling procedure, we analytically derive the first four moments of the conditional average variance, and sample it from the moment-matched shifted lognormal approximation. For the second sampling procedure, we approximate the conditional terminal price as a constant-elasticity-of-variance (CEV) distribution. Our CEV approximation preserves the martingale condition and precludes arbitrage, which is a key advantage over Islah’s approximation used in most SABR simulation schemes in the literature. Then, we adopt the exact sampling method of the CEV distribution based on the shifted-Poisson-mixture Gamma random variable. Our enhanced procedures avoid the tedious Laplace inversion algorithm for sampling integrated variance and non-efficient inverse transform sampling of the forward price in some of the earlier simulation schemes. Numerical results demonstrate our simulation scheme to be highly efficient, accurate, and reliable.
@@
SABR model, Monte Carlo simulation, constant elasticity of variance process, shifted-Poisson-mixture gamma distribution
journal: arXiv.org
@@ 1 Introduction
The stochastic-alpha-beta-rho (SABR) model proposed by Hagan et al. (2002) has been widely adopted in option pricing due to its ability to capture volatility smile or skew using a few parameters, and exhibiting consistency in revealing the dynamic behavior between price and smile. On one hand, it is common in equity options that the implied volatility of out-of-money options are generally higher than that of the in-the-money counterparts, known as the volatility skew. On the other hand, the relatively symmetric volatility smile is common in the foreign exchange options. The SABR model can capture the volatility skew and volatility smile in the respective markets. In addition, the SABR model exhibits consistency in the dynamic behavior of volatility smile or skew, where an increase of the price of the underlying asset leads to shifting of the volatility smile or skew to a higher price (Hagan et al., 2002). The SABR model can generate such co-movements correctly while the local volatility model fails to do so (Derman and Kani, 1998; Dupire, 1997).
@@
Another stream of research, to which this paper belongs, is developing efficient simulation algorithms of the SABR model. With growing popularity of the SABR model, there has been demand for pricing path-dependent derivatives using the model. Since analytic expansion approaches are limited to pricing the European options, the Monte Carlo method becomes a natural choice. In particular, the research on the SABR simulation has been stimulated by less than perfect performance of the Euler and Milstein time-discretization schemes. Since the SABR model imposes an absorbing boundary condition at the origin, the time discretization scheme should be more carefully implemented; naive truncation at the origin leads to large deviation from the original SABR model. Typically, it is necessary to use a small time discretization step to decrease the discretization errors due to its low order of convergence, thus suffering from heavy computation load.
@@
Several simulation algorithms (Chen et al., 2012; Cai et al., 2017; Leitao et al., 2017a, b; Cui et al., 2018; Grzelak et al., 2019; Kyriakou et al., 2023) have been proposed for efficient simulation of the SABR model. Simulating the SABR process over a relatively large time step requires sequentially sampling terminal volatility, average variance, and terminal price at the next simulation time point. Since the last two sampling procedures are challenging, existing methods are differentiated with reference to different choices of the algorithms for these two steps. These earlier simulation algorithms have achieved some success in that they overcome the issue of the time-discretization method. However, there are still many gaps to be filled in order to make SABR simulation more efficient.
@@
Regarding the average variance conditional on the terminal volatility, Chen et al. (2012) sampled the quantity through the mean-and-variance matched lognormal (LN) variable. While the sampling procedures are fast, the time step cannot be large since the derived mean and variance are valid in small-time limit only. For more accurate sampling, researchers used various version of the transform of the average variance. Leitao et al. (2017a, b) derived the approximate Fourier transform of the unconditional average variance recursively, while Cai et al. (2017) adopted the analytically available Laplace transform of the reciprocal of the conditional average variance (Matsumoto and Yor, 2005). However, these approaches are cumbersome and time-consuming due to the tedious numerical inverse transform for constructing the cumulative distribution function (CDF) and iterative root-finding for inverse transform sampling. Cui et al. (2021) approximated the volatility dynamics with the continuous-time Markov chain (CTMC) over volatility grids. Their approach is also not immune from heavy computation and complicated implementation. Kyriakou et al. (2023) used moment-matched Pearson family of distributions for efficient sampling. For such purpose, they rely on numerical evaluation of the four moments from the conditional Laplace transform.
@@
Sampling the terminal price conditional on the terminal variance and average variance is perhaps more challenging. Unlike the other stochastic volatility models that are based on the geometric Brownian motion (BM), the constant-elasticity-of-variance (CEV) feature of the SABR model does not admit an analytically tractable form of the conditional price distribution. As a result, the distribution has to be approximated by an analytic form first even before the corresponding sampling algorithm is considered. For the approximation of the conditional price distribution, almost all existing algorithms (Chen et al., 2012; Cai et al., 2017; Leitao et al., 2017a, b; Cui et al., 2018; Grzelak et al., 2019; Kyriakou et al., 2023) adopted Islah (2009)’s noncentral chi-squared (NCX2) distribution approximation unquestionably. However, the failure of the martingale property in this approximation has not been critically examined among these papers. Leitao et al. (2017a, b) applied an ad-hoc correction on the simulation prices to enforce the martingale property. Therefore, the possibility of better approximation of the conditional price distribution that preserves the martingale should be explored.
@@
The lack of fast sampling algorithm for the approximated distribution is the last hurdle in constructing an efficient SABR simulation algorithm. The inverse transform sampling with numerical root-finding is accurate but tediously slow. The mean-and-variance-match quadratic Gaussian sampling has been adopted in Chen et al. (2012), Leitao et al. (2017a, b), and Cui et al. (2021) to speed up sampling. However, it is limited to small time step due to the nature of approximation. Also, it cannot be used efficiently when the price is close to zero due to the absorbing boundary condition.
@@
This paper proposes an accurate, efficient, and reliable SABR simulation algorithm, filling the gaps in all steps of the simulation procedure. The contributions of this paper are three-fold. Firstly, we derive the first four conditional moments of the conditional average variance analytically. We then use the shifted lognormal (SLN) distribution matching the moments for fast sampling of the average variance. As we use the moments that are higher in order and exact for any arbitrary time step, the validity of our algorithm can be extended to larger time step. Secondly, we present an alternative approximation of the conditional price distribution based on the CEV process. Our approximation is superior to the widely adopted Islah (2009)’s approximation because it preserves the martingale property by construction. Lastly, we employ the algorithms of Makarov and Glew (2010) and Kang (2014) that are capable of sampling the CEV price process for an arbitrary time step. Their algorithms are fast since the CEV price can be expressed by a compound random variable composed of three elementary (i.e., gamma-Poison-gamma) random variables.
@@
This paper is organized as follows. Section 2 presents the formulation of the SABR model and reviews some of its analytic properties. Section 3 describes our SLN approximation for sampling the average variance. In Section 4, we present the CEV distribution approximation of the conditional forward price, and an exact sampling algorithm for the CEV distribution via the shifted Poisson-mixture Gamma distribution. In Sections 3 and 4, our simulation algorithm is compared to those in the existing literature. In particular, we argue that the CEV distribution is a better alternative to the widely adopted Islah (2009)’s approximation as it preserves the martingale condition. Section 5 presents our comprehensive numerical experiments that serve to assess accuracy, efficiency and reliability of different simulation schemes and analytic approximation methods, focusing on the more challenging choices of parameters, such as large time-to-maturity. Conclusive remarks are summarized in Section 6.
2 Formulation of the SABR model and its special cases
@@
In this section, we introduce the SABR model and review some of its analytic properties. The two steps of the SABR simulation are outlined. Specifically, we discuss some special cases where the SABR simulation becomes straightforward.
@@ 2.1 SABR model
The governing stochastic differential equations (SDE) for the SABR volatility model (Hagan et al., 2002) are given by
$\frac{\mathrm{d}F_{t}}{F_{t}^{\beta}}=\sigma_{t}\,\mathrm{d}W_{t}\quad\text{and}\quad\frac{\mathrm{d}\sigma_{t}}{\sigma_{t}}=\nu\,\mathrm{d}Z_{t},\quad(0\leq\beta\leq 1),$ (1)
where $F_{t}$ and $\sigma_{t}$ are the stochastic processes for the forward price and volatility, respectively, $\nu$ is the vol-of-vol, $\beta$ is the elasticity of variance parameter, and $W_{t}$ and $Z_{t}$ are the standard BMs that are correlated with correlation coefficient $\rho$. We also define two variables for later use:
@@
2.2 Average variance, conditional expectation, and two-step simulation procedures
Regarding the simulation step, without loss of generality, we sample the volatility and forward price at time $t+h$ (i.e., $\sigma_{t+h}$ and $F_{t+h}$) for some time step $h$ given time $t$. Accordingly, the filtration up to time $t$ (e.g., $\sigma_{t}$ and $F_{t}$) is implicitly assumed.
Since the volatility process in Eq. (1) follows a geometric BM, $\sigma_{t}=\sigma_{0}\exp\left(\nu Z_{t}-\nu^{2}t/2\right)$, simulating $\sigma_{t}$ from time $t$ to $t+h$ can be easily done by observing
@@ 2.2 Average variance, conditional expectation, and two-step simulation procedures
Here, $\hat{\nu}:=\nu\sqrt{h}$ is the standard deviation of $\ln\sigma_{t}$ over the time step $h$. This quantity is used as a measure of small-time limit.
Like the simulation schemes of other SV models, integrated variance conditional on initial and terminal volatility values plays an important role in the SABR model simulation. Given the initial and terminal volatility (i.e., $\sigma_{t}$ and $\sigma_{t+h}$), we define the conditional average variance between $t$ and $t+h$ as
@@
Here, we normalize the integral by $\sigma_{t}^{2}h$ in order that $I_{t}^{h}$ becomes a dimensionless quantity in the order of one. Specifically, $I_{t}^{h}$ converges to one in the limit of $\nu\downarrow 0$. From the simulation step in Eq. (2), $I_{t}^{h}$ is equivalently conditioned by $\hat{Z}$ as
@@
For notational simplicity, we omit the dependence on $\sigma_{t+h}$ or $\hat{Z}$ and simply write $I_{t}^{h}$.
Almost all SABR simulation algorithms are required to perform the following two steps, given the filtration up to $t$, to sufficient accuracy:
@@
As both simulation steps pose challenges, the previous studies on the topic have proposed numerous algorithms for the two steps. This paper also aims to innovate the two steps, details are presented in Sections 3 and 4, respectively.
Regarding Step 2, let us also define the conditional expectation of $F_{t+h}$:
@@
which is essential for the discussion of the simulation algorithms of Step 2. We also omit the dependence on $\sigma_{t+h}$ and $I_{t}^{h}$, and simply write $\bar{F}_{t}^{h}$.
@@ 2.3 Special cases
Before introducing our simulation algorithm, we review several special cases of the SABR model. In these cases, Step 2 is relatively easier while Step 1 remains challenging. Reviewing these cases also helps building intuitions for our new algorithm.
@@
$F_{T}\,|\,\sigma_{T},I_{0}^{T}\;\sim\;\mathcal{N}\left(\bar{F}_{0}^{T},\rho_{\ast}^{2}\sigma_{0}^{2}TI_{0}^{T}\right)$
@@
Therefore, Step 2 becomes trivial as long as $I_{t}^{h}$ can be sampled accurately in Step 1. Choi et al. (2019) have also found an exact closed-form expression for sampling $F_{T}$ without separating Steps 1 and 2. Moreover, note that $F_{T}$, as well as $\bar{F}_{0}^{T}$ can be negative in this special case as the origin is not a boundary. The normal SABR model is different from the $\beta\downarrow 0$ limit of the SABR model, where the origin is an absorbing boundary. In the view of the above, the SABR model under $\beta=0$ is not a concern of our study.
Lognormal ($\beta=1$) SABR model: When $\beta=1$, conditional on $\sigma_{T}$ and $I_{0}^{T}$, integrating $\ln F_{t}$ from $t=0$ to $T$ yields that the terminal forward price $F_{T}$ follows the LN distribution with mean $\bar{F}_{0}^{T}$ and log variance $\rho_{\ast}^{2}\sigma_{0}^{2}TI_{0}^{T}$:
@@
where $X$ is a standard normal variate independent of $\sigma_{T}$ and $I_{0}^{T}$. The conditional expectation $\bar{F}_{0}^{T}$ is given by
@@
This special case provides an important observation, which derives an intuition for our CEV approximation when $10<\beta<$ 1 (to be discussed Section 4).
@@
The joint distribution of $\sigma_{T}$ and $I_{0}^{T}$ satisfies
@@
Proof.
It follows from the law of total expectation:
$E(\bar{F}_{0}^{T})=E\left(E\left(F_{T}|\sigma_{T},I_{0}^{T}\right)\right)=E(F_{T})=F_{0}.\qed$
Zero vol-of-vol ($\nu=0$) SABR model: When $10<\beta<$ 1, the SABR model is visualized as the stochastic volatility version of the CEV model:
$\frac{\mathrm{d}F_{t}}{F_{t}^{\beta}}=\sigma_{0}\,\mathrm{d}W_{t},$ (8)
with the absorbing boundary condition at the origin.
@@
We define the distribution of $F_{T}$ resulting from the CEV dynamics in Eq. (8) as the CEV distribution,
$F_{T}\sim\mathcal{CEV}_{\beta}(F_{0},\sigma_{0}^{2}T),$
parameterized by the elasticity-of-variance parameter $\beta$, mean $F_{0}$ [^1] If $F_{T}\sim\mathcal{CEV}_{\beta}(F_{0},\sigma_{0}^{2}T)$, then $E(F_{T})=F_{0}$ regardless of the values of $\beta$ and $\sigma_{0}^{2}T$., and variance $\sigma_{0}^{2}T$. The probability distribution function (PDF) and CDF of the CEV distribution will be shown later in Proposition 5.
Therefore, in the zero vol-of-vol ($\nu\downarrow 0$) limit, $F_{T}$ from the SABR model (unconditionally) follows $\mathcal{CEV}_{\beta}(F_{0},\sigma_{0}^{2}T)$. Step 2 is reduced to sampling $F_{T}$ from the CEV distribution. However, this is not a trivial task as we discuss later in Section 4.
@@
Basically, $TI_{0}^{T}$ plays the role of stochastic time clock of the CEV model. As in the $\nu=0$ case, Step 2 is reduced to sampling $F_{T}$ from the CEV distribution in this case. Unlike the $\beta=1$ case, $F_{T}$ is conditioned by $I_{0}^{T}$ only since conditional expectation is always $F_{0}$ without depending on $\sigma_{T}$. Thus, $I_{0}^{T}$ should be understood as the unconditional average variance in this context. Therefore, the European option price can be simplified to one dimensional integral of the CEV option price over the (unconditional) distribution of $I_{0}^{T}$. Choi and Wu (2021b) have approximated the integral using the Gaussian quadrature together with the LN distribution fitted to the (unconditional) mean and variance of $I_{0}^{T}$.
Completing the discussion on the four special cases, we are left with the most general case with $10<\beta<$ 1 and $\rho\neq 0$, under which the conditional distribution of $F_{T}$ is not analytically tractable. One has to resort to an approximation. The approximation should be reduced to the special cases discussed above under various continuous limits.

### v2

# Efficient and accurate simulation of the stochastic-alpha-beta-rho model

@@ Abstract
We propose an efficient, accurate and reliable simulation scheme for the stochastic-alpha-beta-rho (SABR) model. The two challenges of the SABR simulation lie in sampling (i) integrated variance conditional on terminal volatility and (ii) terminal forward price conditional on terminal volatility and integrated variance. For the first sampling procedure, we sample the conditional integrated variance using the moment-matched shifted lognormal approximation. For the second sampling procedure, we approximate the conditional terminal forward price as a constant-elasticity-of-variance (CEV) distribution. Our CEV approximation preserves the martingale condition and precludes arbitrage, which is a key advantage over Islah’s approximation used in most SABR simulation schemes in the literature. We then adopt the exact sampling method of the CEV distribution based on the shifted-Poisson mixture Gamma random variable. Our enhanced procedures avoid the tedious Laplace inversion algorithm for sampling integrated variance and non-efficient inverse transform sampling of the forward price in some of the earlier simulation schemes. Numerical results demonstrate our simulation scheme to be highly efficient, accurate, and reliable.
@@
SABR model, constant elasticity of variance process, simulation, shifted-Poisson-mixture Gamma distribution
journal: European Journal of Operational Research
@@ 1 Introduction
The stochastic-alpha-beta-rho (SABR) model proposed by Hagan et al. (2002) has been widely adopted in option pricing due to its ability to capture volatility smile or skew using a few parameters, and exhibiting consistency in revealing the dynamic behavior between asset price and volatility smile. On one hand, it is common in equity options that the implied volatility of out-of-money options is generally higher than that of the in-the-money counterparts, known as the volatility skew. On the other hand, the relatively symmetric volatility smile is common in the foreign exchange options. The SABR model can capture the volatility skew and volatility smile in the respective markets. In addition, the SABR model exhibits consistency in the dynamic behavior of volatility smile or skew, where an increase of the price of the underlying asset leads to shifting of the volatility smile or skew to a higher price (Hagan et al., 2002). The SABR model can generate such co-movements correctly while the local volatility model fails to do so (Derman and Kani, 1998; Dupire, 1997).
@@
This paper follows the other stream of research, which is the development of efficient simulation algorithms for pricing options under the SABR model. With growing popularity of the SABR model, there has been demand for pricing path-dependent derivatives using the model. Since analytic approximation approaches are limited to pricing European options, the Monte Carlo method becomes a better choice. In particular, the research on the SABR simulation started with the less than perfect performance of the Euler and Milstein time-discretization schemes. Since the SABR model imposes an absorbing boundary condition at the origin, the numerical boundary condition in the time discretization schemes have to be more carefully implemented. It has been well known that naive truncation at the origin may lead to large deviation from the original SABR model. Typically, it is necessary to use a small time discretization step to decrease the discretization errors due to its low order of convergence, thus causing heavy computation load.
@@
Several simulation algorithms (Chen et al., 2012; Cai et al., 2017; Leitao et al., 2017a, b; Cui et al., 2018; Grzelak et al., 2019; Kyriakou et al., 2023) have been proposed for the simulation of the SABR model. Simulation of the SABR model over one time step requires sequential sampling procedures of the terminal volatility, average variance, and terminal forward price at the next simulation time point. Since the last two sampling procedures are challenging, existing methods distinguish among themselves by adopting different choices of the algorithms for these two steps. These earlier simulation algorithms have achieved some successes to overcome the difficulty of imposing the absorbing boundary condition in the time-discretization methods. This paper aims to provide more efficient, accurate and reliable simulation scheme that competes favorably with the existing simulation schemes in the literature.
@@
Regarding the sampling of the average variance conditional on the terminal volatility, Chen et al. (2012) sampled this quantity through the mean-and-variance matched lognormal (LN) variable. While the sampling procedures are fast, the time step cannot be large since the derived mean and variance are valid in small-time limit only. For more accurate sampling, later works employed various versions of the transform of the average variance. Leitao et al. (2017a, b) derived the approximate Fourier transform of the unconditional average variance recursively, while Cai et al. (2017) used the analytic formula of the Laplace transform of the reciprocal of the conditional average variance (Matsumoto and Yor, 2005). However, these approaches are cumbersome and time-consuming since the tedious numerical inverse transform procedures are required to construct the cumulative distribution function (CDF) and implement the iterative root-finding for the inverse transform sampling. As an alternative approach, Cui et al. (2021) approximated the volatility dynamics with the continuous-time Markov chain (CTMC) over volatility grids. Unfortunately, their approach does not immune from heavy computation and cumbersome implementation. In another approach, Kyriakou et al. (2023) used the moment-matched Pearson family of distributions for efficient sampling. For such purpose, they rely on numerical evaluation of the four order moments from the Laplace transform of conditional average variance.
@@
Sampling the terminal forward price conditional on the terminal variance and average variance is much more challenging. Unlike the other stochastic volatility models that are based on the geometric Brownian motion (BM), the constant-elasticity-of-variance (CEV) feature of the SABR model does not admit an analytically tractable form of the conditional terminal forward price distribution. As a result, the price distribution has to be approximated in some analytic form first before the corresponding sampling algorithm is considered. For the approximation of the conditional terminal forward price distribution, most existing algorithms (Chen et al., 2012; Cai et al., 2017; Leitao et al., 2017a, b; Cui et al., 2018; Grzelak et al., 2019; Kyriakou et al., 2023) invariably adopted Islah (2009)’s noncentral chi-squared (NCX2) distribution approximation. However, the failure of the martingale property of the conditional forward price in this approximation has not been critically examined among these earlier schemes. Leitao et al. (2017a, b) applied an ad-hoc correction on the simulation prices to enforce the martingale property. Our paper proposes the CEV approximation of the conditional terminal forward price distribution, which is considered to be better since it preserves the martingale property and provides more efficient simulation procedures.
@@
Indeed, the lack of fast sampling algorithm for the approximated conditional forward price distribution is the last challenge in constructing an efficient SABR simulation algorithm. The inverse transform sampling with numerical root-finding procedure is accurate but tediously slow. The mean-and-variance-match quadratic Gaussian sampling has been adopted in Chen et al. (2012), Leitao et al. (2017a, b), and Cui et al. (2021) to speed up sampling. However, it is limited to small time step due to the nature of approximation. Also, it cannot be used efficiently when the price is close to zero due to the absorbing boundary condition.
@@
This paper proposes an accurate, efficient, and reliable SABR simulation algorithm, filling the gaps in all steps of the simulation procedure. The contributions of this paper are three-fold. Firstly, we derive the first four moments of the conditional average variance analytically. We then use the shifted lognormal (SLN) distribution obtained from matching the moments for fast sampling of the average variance. Since we match the moments that are in higher order and exact for any arbitrary time step, good accuracy of our SLN distribution approximation can be maintained at larger time step. Secondly, we present an alternative approximation of the conditional forward price distribution based on the CEV process. Our approximation shows better accuracy when compared with the widely adopted Islah’s approximation (Islah, 2009) since it preserves the martingale property of the conditional forward price by its construction. Lastly, we employ the algorithms of Makarov and Glew (2010) and Kang (2014) that are capable of sampling the CEV process exactly. Their algorithms are fast since the CEV distribution can be expressed by a compound random variable composed of three elementary (namely, gamma-Poisson-gamma) random variables.
@@
This paper is organized as follows. Section 2 presents the formulation of the SABR model and reviews some of its analytic properties. Section 3 describes our SLN approximation for sampling the average variance and presents its performance in numerical accuracy. In Section 4, we present the CEV distribution approximation of the conditional forward price, and an exact sampling algorithm for the CEV distribution via the shifted Poisson-mixture Gamma distribution. In particular, we argue that our CEV distribution approximation of the conditional forward price is better than the widely adopted Islah’s approximation as it preserves the martingale condition. Section 5 presents our comprehensive numerical experiments that serve to assess accuracy, efficiency and reliability of different simulation schemes and analytic approximation methods, displaying numerical comparisons under the more challenging choices of parameters, such as large time-to-maturity. Comparison of accuracy performance of our CEV approximation of the forward price with that of the Islah’s approximation is carefully examined. Conclusive remarks are summarized in Section 6.
2 Formulation of the SABR model and its two-step simulation
@@
In this section, we introduce the SABR model and review some of its analytic properties under various degenerate cases. The two steps of the SABR simulation are outlined. Specifically, we discuss some special cases where the SABR simulation becomes straightforward.
@@ 2.1 SABR model
The governing stochastic differential equations (SDEs) for the SABR volatility model (Hagan et al., 2002) are given by
$\frac{\mathrm{d}F_{t}}{F_{t}^{\beta}}=\sigma_{t}\,\mathrm{d}W_{t}\quad\text{and}\quad\frac{\mathrm{d}\sigma_{t}}{\sigma_{t}}=\nu\,\mathrm{d}Z_{t},\quad 0\leq\beta\leq 1,$ (1)
where $F_{t}$ and $\sigma_{t}$ are the stochastic processes for the forward price and volatility, respectively, $\nu$ is the vol-of-vol, $\beta$ is the elasticity of variance parameter, and $W_{t}$ and $Z_{t}$ are the standard Brownian motions (BMs) that are correlated with correlation coefficient $\rho$. We also define two parameters for notational simplicity:
@@
2.2 Average variance and two-step simulation procedures
Regarding the simulation over one time step with size $h$, we sample the volatility and forward price at time $t+h$ ($\sigma_{t+h}$ and $F_{t+h}$) for discrete time step $h$ at given time $t$. Accordingly, the filtration up to time $t$ (say, $\sigma_{t}$ and $F_{t}$) is implicitly assumed. Since the volatility process $\sigma_{t}$ in Eq. (1) follows a geometric BM, we have $\sigma_{t}=\sigma_{0}\exp\left(\nu Z_{t}-\nu^{2}t/2\right).$ Simulating $\sigma_{t}$ from time $t$ to $t+h$ can be easily performed via
@@ 2.2 Average variance and two-step simulation procedures
Here, $\hat{\nu}:=\nu\sqrt{h}$ is the standard deviation of $\ln\sigma_{t}$ over the time step $h$. The quantity $\hat{\nu}$ is used as a measure of small-time limit.
Like the simulation schemes in other stochastic volatility models, the average variance conditional on initial and terminal volatility values plays an important role in the SABR model simulation. Given the initial and terminal volatility (namely, $\sigma_{t}$ and $\sigma_{t+h}$), we define the dimensionless conditional average variance between $t$ and $t+h$ as
@@
Here, we normalize the integrated variance by $\sigma_{t}^{2}h$ in order that $I_{t}^{h}$ becomes a dimensionless quantity of order one. Specifically, $I_{t}^{h}$ converges to one in the limit of $\nu\downarrow 0$. From the simulation step in Eq. (2), $I_{t}^{h}$ is equivalently conditioned by $\hat{Z}$ as
@@
For notational simplicity, we simply write $I_{t}^{h}$ by omitting the dependence on $\sigma_{t+h}$ or $\hat{Z}$.
Almost all SABR simulation algorithms strive to perform the following two steps of simulation to achieve efficiency and high accuracy, given the filtration up to $t$:
@@
This paper aims to innovate these two steps of simulation over existing algorithms, details are presented in Sections 3 and 4, respectively.
Regarding Step 2, we define the conditional expectation of $F_{t+h}$:
@@
which is essential for the discussion of the simulation algorithms of Step 2. For notational simplicity, we omit the dependence on $\sigma_{t+h}$ and $I_{t}^{h}$, and simply write $\bar{F}_{t}^{h}$.
@@ 2.3 Special cases
Before introducing our simulation algorithm, we review several special cases of the SABR model, which helps building intuition for our new algorithm.
@@
$F_{T}\,|\,\sigma_{T},I_{0}^{T}\;\sim\;\mathcal{N}\left(\bar{F}_{0}^{T},\rho_{\ast}^{2}\sigma_{0}^{2}TI_{0}^{T}\right),$
@@
Therefore, Step 2 becomes trivial as long as $I_{t}^{h}$ can be sampled accurately in Step 1. Choi et al. (2019) managed to find an exact closed-form expression for sampling $F_{T}$ without separating Steps 1 and 2. Moreover, note that $F_{T}$, as well as $\bar{F}_{0}^{T}$ can be negative under normal distribution as the origin is not a boundary. Accordingly, the normal SABR model is different from the $\beta\downarrow 0$ limit of the SABR model. In the later case, the origin is an absorbing boundary. In view of these subtleties, we discard further discussion of the SABR model under $\beta=0$.
Lognormal ($\beta=1$) SABR model: When $\beta=1$, conditional on $\sigma_{T}$ and $I_{0}^{T}$, integrating $\ln F_{t}$ from $t=0$ to $T$ reveals that the terminal forward price $F_{T}$ follows the LN distribution with mean $\bar{F}_{0}^{T}$ and log variance $\rho_{\ast}^{2}\sigma_{0}^{2}TI_{0}^{T}$:
@@
where $X$ is a standard normal variate independent of $\sigma_{T}$ and $I_{0}^{T}$. The conditional expectation $\bar{F}_{0}^{T}$ is explicitly given by
@@
This special case of $\beta=1$ provides the important intuition for our CEV approximation of the forward price under $0<\beta<1$ (to be discussed in Section 4).
@@
From the law of total expectation, it follows that
$E(\bar{F}_{0}^{T})=E\left(E\left(F_{T}|\sigma_{T},I_{0}^{T}\right)\right)=E(F_{T})=F_{0},$
and the joint distribution of $\sigma_{T}$ and $I_{0}^{T}$ satisfies
@@
Zero vol-of-vol ($\nu=0$) SABR model: For $0<\beta<1$, the SABR model is now reduced to the CEV model:
$\frac{\mathrm{d}F_{t}}{F_{t}^{\beta}}=\sigma_{0}\,\mathrm{d}W_{t},$ (8a)
with the absorbing boundary condition at the origin. Here, $\sigma_{0}$ is constant. Indeed, the SABR model can be visualized as the stochastic volatility version of the CEV model.
@@
The CEV distribution is defined as the distribution of $F_{T}$ resulting from the CEV dynamics as depicted in Eq. (8a), which is denoted by
$F_{T}\sim\mathcal{CEV}_{\beta}(F_{0},\sigma_{0}^{2}T).$ (8b)
It is parameterized by the elasticity-of-variance parameter $\beta$, mean $F_{0}$, and variance $\sigma_{0}^{2}T$. The cumulative distribution function (CDF) and probability density function (PDF) of the CEV distribution are presented respectively in Eqs. (22a) and  (22b) later.
If $F_{T}\sim\mathcal{CEV}_{\beta}(F_{0},\sigma_{0}^{2}T)$, then $E(F_{T})=F_{0}$ regardless of the values of $\beta$ and $\sigma_{0}^{2}T$. In the zero vol-of-vol ($\nu\downarrow 0$) limit, $F_{T}$ from the SABR model follows $\mathcal{CEV}_{\beta}(F_{0},\sigma_{0}^{2}T)$ unconditionally. Step 2 is reduced to sampling $F_{T}$ from the CEV distribution.
@@
Basically, $TI_{0}^{T}$ plays the role of stochastic time clock of the CEV model. Like the $\nu=0$ case, Step 2 is reduced to sampling $F_{T}$ from the CEV distribution under $\rho=0$. Unlike the $\beta=1$ case, $F_{T}$ is only conditioned by $I_{0}^{T}$ since conditional expectation is always $F_{0}$ without depending on $\sigma_{T}$. Thus, $I_{0}^{T}$ should be understood as the unconditional average variance in this context. Therefore, the European option price can be expressed as one dimensional integral of the CEV option price over the (unconditional) distribution of $I_{0}^{T}$. As a remark, Choi and Wu (2021b) approximated the integral using the Gaussian quadrature together with the LN distribution fitted to the (unconditional) mean and variance of $I_{0}^{T}$.
Besides the above four special cases, one has to face with the general case with $0<\beta<1$ and $\rho\neq 0$, under which the conditional distribution of $F_{T}$ is not analytically tractable. The success of a simulation algorithm for the SABR model amounts to an ingenious choice chosen for the approximation of the conditional distribution of $F_{T}$.